OpenStudy (anonymous):

Devaughn's age is two times Sydney's age. The sum of their ages is . What is Sydney's age?

OpenStudy (lgbasallote):

let d= devaughn's age s = sydney's age the sum of their ages is what..?? o.O

OpenStudy (anonymous):

am i suppose to divde?

OpenStudy (anonymous):

you need to set it up as: d=devaughn's age s=sydney's age you have 2 unknowns but you also have 2 equations: d+s= whatever d=2s by substituting you get 3s=whatever, so you can solve for sydney's age, then use it to get devaughn's age

OpenStudy (lgbasallote):

well you can just stop at sydney's age haha

OpenStudy (anonymous):

or that

OpenStudy (anonymous):

With d=3s..... won't it be d=3(66)?

OpenStudy (lgbasallote):

hmmm..why dont you tell us what the sum of their ages is first..before you ask us to verify your answers..good idea? :D

OpenStudy (anonymous):

you never told us their combined age, we can't check it

OpenStudy (lgbasallote):

though seeing the equation right now...i dont think devaughn should be around 180 yrs old...

OpenStudy (anonymous):

"Devaughn's age (d) is two times Sydney's age (s). The sum of their ages is T. What is Sydney's age? " So from the first sentence, d =2s. From the second sentence, d + s = T. Now solve for s. Substitute the first equation into the second, yeilding 2s + s = T. Solve for s.